OpenStudy (anonymous):

simplify: 2√18a^2x^5

OpenStudy (tkhunny):

Sadly, we really can't tell what is under the radical. Have you considered finding perfect squares under the radical an pulling them out? Like This sqrt(24) = sqrt(4*6) = sqrt(4)*sqrt(6) = 2*sqrt(6)

OpenStudy (anonymous):

Eveerything from the 18 is under the radical , do you mind explaining where you got those numbers from? I'm sorry math is just my worst subject

OpenStudy (tkhunny):

I made up an example. You try it with 18. You need to get good at "Prime Factorization". Have you done it?

OpenStudy (anonymous):

Ohh okay ill try it now

OpenStudy (triciaal):